OpenStudy (anonymous):

Two cars collide, car one has a mass of 1000Kg and is travelling at 10m/s east to west. Car two has a mass of 1500Kg and is travelling at an unknown velocity west to east. After collision they are both travelling at 2m/s west to east. Calculate the velocity of car two before the collision.

OpenStudy (anonymous):

I have been doing this but my answer looks wrong. P=MV p=1000*20=20000. rearrange p=mv to make m=v/p so m=40/20000 = 0.002 this is obviously wrong but I don't know why?

OpenStudy (anonymous):

do you have the ans.........

OpenStudy (anonymous):

I figured it out im stupid

OpenStudy (anonymous):

I rearranged the formula wrong. Im tired lol

OpenStudy (anonymous):

For anyone who is looking the answer was 500KG

OpenStudy (anonymous):

i thought you wanted the velocity before the collision of the other car

OpenStudy (anonymous):

m= mass, v = velocity m1V1initial+m2V2initial = m1V1final +m2V2final think the answer is -3.3 m/s before they collide.

OpenStudy (anonymous):

i meant 3.3 m/s well hope i get a medal for trying! LoL peace!

OpenStudy (anonymous):

Oh thanks it seems the question I asked originally was not the one I was working out lol. Very tired